OpenStudy (anonymous):

Alfredo has forgotten the five-digit code to the lock on his suitcase. Alfredo remembered that the first digit is 0 and the last digit is either 8 or 4. The middle digits could be any number from 0 through 9 inclusive. What is the greatest number of five-digit codes Alfredo might have to try in order to be sure to open his suitcase? 1,458 32 29 2,000

OpenStudy (anonymous):

2 times (max # for a position)^3

OpenStudy (anonymous):

go with the last one you have 10 choices for the second third and 4th digit, 2 for the fifth counting principle says to multiply all these together \[10\times 10\times 10\times 2\]

OpenStudy (anonymous):

So, what is the max # for a position? Look at position 2 and it goes from 0 to 10.

OpenStudy (anonymous):

max # for position is 10?

OpenStudy (anonymous):

Yes! Good work! so, 2 x 10^3

OpenStudy (anonymous):

2000?

OpenStudy (anonymous):

You got it, my friend!

OpenStudy (anonymous):

thx :D

OpenStudy (anonymous):

You're welcome. Satellite, thx for the recognition!